OpenStudy (anonymous):

Which number line represents the solution of x2 - 16 ≤ 0?

OpenStudy (anonymous):

x^2<=16 x^2<=(+- 4)^2

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

Now, HINT: If x^2<=(+- a)^2 then -a<=x<=a

OpenStudy (anonymous):

Does it help?